Exhibit 24-10

Quantity Sold

Price

(units)

Total Cost

$10

10

$80

9

20

100

8

30

130

7

40

170

6

50

230

5

60

300

4

70

380

Refer to Exhibit 24-9. A single-price monopolist earns a total profit of __________ when it produces and sells 40 units of its good.

Group of answer choices

$80.

$110.

$30.

$61.

$49.
Answer #1

Option B.

At 40 units of the output produced and sold, the total revenue of the firm is equal to = 40 * 7 = $280. Total cost to produce 40 units of output = $170, thus, profits of the firm = Total Revenue - Total Cost = 280 - 170 = $110.